### v1

@@ Appendix B Auxiliary Proofs
Lemma 16 (Bounding Forgetting Using the Training Loss).
In a realizable setting (Assumption 1), the iterates of Scheme 1 under a random task ordering $\tau$ (with or without replacement) hold $\forall k\geq 1$,
@@
As discussed in Section 3.1, Scheme 2 governs the updates of the iterates $\mathbf{w}_{t}\in\mathbb{R}^{d}$. Under Assumption 1, we define the orthogonal projection as ${\mathbf{P}}_{\tau(t)}\triangleq\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}$, revealing a recursive form:
@@
[Assumption 1]delimited-[]Assumption $\end{subarray}\right]$ $=\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\mathbf{w}_{\star}+\left(\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\right)\mathbf{w}_{t-1}=(\mathbf{I}_{d}-{\mathbf{P}}_{\tau\left({t}\right)})\mathbf{w}_{\star}+{\mathbf{P}}_{\tau\left({t}\right)}\mathbf{w}_{t-1}$
@@
$\frac{R^{2}}{k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left({\mathbf{w}_{t-1}-\mathbf{w}_{\star}}\right)\right\|^{2}$ $=\frac{R^{2}}{k}\sum_{t=1}^{k}\left(\mathbb{E}_{\tau}\left\|\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right\|^{2}-\mathbb{E}_{\tau}\left\|{\mathbf{P}}_{\tau\left(t\right)}\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\right)$
@@
Proposition 17 (Bounding The Training Loss Using Forgetting in Without-Replacement Orderings).
@@
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,,$
@@
$\mathcal{L}\left(\mathbf{w}_{k}\right)$ $=\frac{1}{2T}\sum_{m=1}^{T}\left\|\mathbf{w}_{k}-{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2},\quad F_{\tau}\left(k\right)=\frac{1}{2k}\sum_{t=1}^{k}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(t\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,.$
@@
For the POCS case, simply replace $\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$ with $\left\|\mathbf{w}_{k}-{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2}$.
  • Recall Lemma 6.Consider any realizable task collection such that $\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m},\forall m\in\left[{T}\right]$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
@@
    Finally, the convexity of $f_{m}$ is immediate since $\nabla_{\mathbf{w}}^{2}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\succeq\mathbf{0}$.
@@
  • Recall Theorem 7.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 3$ iterations are upper bounded as,
@@
    By defining $\mathbf{z}_{t}=\mathbf{w}_{t}-\mathbf{w}_{\star}$ and exploiting the recursive form of Eq. 4 from the proof of Lemma 16, we obtain $\mathbf{z}_{t}={\mathbf{P}}_{\tau(t)}\mathbf{z}_{t-1}$. Note that $\mathbf{z}_{0}=\mathbf{0}_{d}-\mathbf{w}_{\star}=-\mathbf{w}_{\star}$.
@@
    This map plays a central role in our analysis and has been studied in similar forms in prior work (Guo et al., 2022). Note that ${\mathbf{P}}_{m}$ is an orthogonal projection, which implies that it is symmetric and idempotent. Thus,
@@
    It follows that $\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]=Q^{t}\left[{\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]}\right]=Q^{t}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]=Q^{t}\left[{\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)^{\top}}\right]=Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]$, where $Q^{t}$ denotes $t$ applications of $Q$. The map $Q$ captures the evolution of the error’s second-moment under Kaczmarz updates, offering a tractable approach to analyzing the algorithm’s convergence. The expected loss at step $t$ is given by
    $\mathbb{E}_{\tau}\left[{\mathcal{L}\left({\mathbf{w}_{t}}\right)}\right]$ $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{w}_{t}-\mathbf{y}_{i}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\left({\mathbf{w}_{t}-\mathbf{w}_{\star}}\right)}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{z}_{t}}\right\|^{2}}\right]$
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\left\|{\mathbf{X}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\mathbf{z}_{t}^{\top}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right)}\right]$
@@
    We are now ready to derive the final bound. From Lemma 24, we have
@@
    Additionally, by Corollary 22, $Q^{k}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]$ is symmetric and positive semidefinite (PSD). We also note that $\frac{1}{T}\mathbf{X}^{\top}\mathbf{X}$ is symmetric PSD. The key insight from Lemma 24, combined with the trace product inequality (Lemma 23), is that it allows the expected loss to be expressed using a polynomial in $Q$. This reformulation simplifies the convergence analysis by reducing it to examining the spectral properties of $Q$. Invoking the trace product inequality, we obtain:
@@
    To clarify, the operator norm of a linear map $H$ is defined as $\left\|{H}\right\|=\sup_{\mathbf{A}\in\mathbb{R}^{d\times d},\left\|{\mathbf{A}}\right\|_{F}=1}\left\|{H\left[{\mathbf{A}}\right]}\right\|_{F}$. The reason for switching from the spectral norm to the Frobenius norm is to enable the use of the spectral mapping theorem to bound the operator norm of $Q^{k-1}\left({I-Q}\right)$, applicable only for inner-product-based norms. We complete the proof by bounding the forgetting using the training loss (Lemma 16). That is,
@@ C.1 Key Properties and Auxiliary Lemmas
Let $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, for all $i\in\left[{T}\right]$, $\mathbf{0}\preccurlyeq{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}$. Meaning $Q\left[{\mathbf{A}}\right]$ is PSD as a convex combination of PSD matrices.
Corollary 21.
@@
For $n=1$, given by Corollaries 20 and 21. For $n>1$, this follows trivially by induction.
Lemma 23 (Trace Product Inequality).
@@
Since $\mathbf{0}\preccurlyeq\mathbf{C}=\mathbf{C}^{\top}$, it has a square symmetric PSD root $\mathbf{C}^{1/2}$. Given that $\mathbf{A},\mathbf{B}$ are symmetric and $\mathbf{A}\preccurlyeq\mathbf{B}$, it follows that $\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}\preccurlyeq\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}$ (from Horn and Johnson, 2012, Theorem 7.7.2.a). Applying the cyclic property of the trace and using the fact that for symmetric matrices ordered in the Löwner sense, their traces are also ordered (Horn and Johnson, 2012, Corollary 7.7.4.d), we obtain
@@
$\left[\begin{subarray}{c}\text{Im}\left({\mathbf{X}_{i}^{+}\mathbf{X}_{i}}\right)\subseteq\text{Im}\left({\mathbf{X}^{+}\mathbf{X}}\right)\end{subarray}\right]$ $=-\frac{1}{T}\sum_{i=1}^{T}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}+\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{V}_{i}{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}\mathbf{V}_{i}^{\top}.$
@@
[Corollary 21]delimited-[]Corollary $\end{subarray}\right]$ $=\operatorname{tr}\left({Q\left[{\mathbf{B}}\right]^{\top}\mathbf{A}}\right)=\operatorname{tr}\left({\mathbf{A}^{\top}Q\left[{\mathbf{B}}\right]}\right)=\langle\mathbf{A},Q\left[{\mathbf{B}}\right]\rangle\,.$
Proposition 27.
@@
From the Rayleigh quotient characterization of eigenvalues, this implies that every eigenvalue $\lambda$ of $Q$ satisfies $0\leq\lambda\leq 1$, i.e., $\sigma(Q)\subset[0,1]\,$.
Lemma 28.
@@
By Proposition 26, $Q$ is self adjoint. Thus, we can apply the spectral mapping theorem to the polynomial $x\mapsto x^{n}\left({1-x}\right)$. The eigenvalues of $Q^{n}\left({I-Q}\right)$ are of the form $\lambda^{n}\left({1-\lambda}\right)$, where $\lambda$ is an eigenvalue of $Q$. From Proposition 27, we know that $\lambda\in\left[0,1\right]$. Using an algebraic property of $\lambda^{n}\left({1-\lambda}\right)$ for $\lambda\in\left[{0,1}\right]$, we conclude that $\lambda^{n}\left({1-\lambda}\right)\in\left[0,\frac{1}{en}\right]$.
Therefore, $\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en}$.
Lemma 29.
@@
We first bound $\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$ using the operator norm bound on $Q$ (Proposition 27):
@@
The proofs in this appendix focus on the properties of forgetting and loss, “translating” them into the language of last-iterate SGD. We then apply our last-iterate results, proved in Appendix E.
D.1 Proof of Theorem 8: A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
  • Recall Theorem 8.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 2$ iterations are bounded as,
@@ D.1 Proof of Theorem 8: A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
    Furthermore, Lemma 6 states that for all $\mathbf{w}\in\mathbb{R}^{d}$,
@@
    will imply the desired result. Indeed, again by Lemma 6, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$. Hence, plugging in $\mathbf{A}=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\Rightarrow\left\|{\mathbf{A}}\right\|=1=\beta$ into Theorem 10, SGD with $\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@
    and therefore $\mathbb{E}\mathcal{L}(\mathbf{w}_{k})\leq\frac{2R^{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}$, which proves the first claim. The second claim follows immediately from Lemma 16, and we are done.
D.2 Proving Theorem 9: Main Result for Without Replacement Orderings
  • Recall Theorem 9.Under a random ordering without replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\in\{2,\dots,T\}$ iterations are both bounded as,
@@ D.2 Proving Theorem 9: Main Result for Without Replacement Orderings
From Lemmas 16 and 6, we have
@@
Combining with Proposition 17, we get,
@@
Recall that from Eq. 5 in the proof of Lemma 16, we have
@@
$\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({t}\right)}}\right){\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right)=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right).$

### v2

@@ Appendix B Auxiliary Proofs
Lemma B.1 (Bounding Forgetting Using the Training Loss).
In a realizable setting (Assumption 2.2), the iterates of Scheme 1 under a random task ordering $\tau$ (with or without replacement) hold $\forall k\geq 1$,
@@
    As discussed in Section 3.1, Scheme 2 governs the updates of the iterates $\mathbf{w}_{t}\in\mathbb{R}^{d}$. Under Assumption 2.2, we define the orthogonal projection as ${\mathbf{P}}_{\tau(t)}\triangleq\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}$, revealing a recursive form:
@@
    [Assumption 2.2]delimited-[]Assumption $\end{subarray}\right]$ $=\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\mathbf{w}_{\star}+\left(\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\right)\mathbf{w}_{t-1}=(\mathbf{I}_{d}-{\mathbf{P}}_{\tau\left({t}\right)})\mathbf{w}_{\star}+{\mathbf{P}}_{\tau\left({t}\right)}\mathbf{w}_{t-1}$
@@
    $\frac{R^{2}}{k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left({\mathbf{w}_{t-1}-\mathbf{w}_{\star}}\right)\right\|^{2}$
    $=\frac{R^{2}}{k}\sum_{t=1}^{k}\left(\mathbb{E}_{\tau}\left\|\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right\|^{2}-\mathbb{E}_{\tau}\left\|{\mathbf{P}}_{\tau\left(t\right)}\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\right)$
@@
$\blacksquare$
Proposition B.2 (Bounding The Training Loss Using Forgetting in Without-Replacement Orderings).
@@
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-\bm{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,,$
@@
$\mathcal{L}\left(\mathbf{w}_{k}\right)$ $=\frac{1}{2T}\sum_{m=1}^{T}\left\|\mathbf{w}_{k}-\bm{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2},\quad F_{\tau}\left(k\right)=\frac{1}{2k}\sum_{t=1}^{k}\left\|\mathbf{w}_{k}-\bm{\Pi}_{\tau\left(t\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,.$
@@
    For the POCS case, simply replace $\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$ with $\left\|\mathbf{w}_{k}-\bm{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2}$. $\blacksquare$
  • Recall Lemma 3.1.Consider any realizable task collection such that $\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m},\forall m\in\left[{T}\right]$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
@@
      Finally, the convexity of $f_{m}$ is immediate since $\nabla_{\mathbf{w}}^{2}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\succeq\mathbf{0}$. $\blacksquare$
@@
  • Recall Theorem 4.1.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 3$ iterations are upper bounded as,
@@
    By defining $\mathbf{z}_{t}=\mathbf{w}_{t}-\mathbf{w}_{\star}$ and exploiting the recursive form of Eq. 3 from the proof of Lemma B.1, we obtain $\mathbf{z}_{t}={\mathbf{P}}_{\tau(t)}\mathbf{z}_{t-1}$. Note that $\mathbf{z}_{0}=\mathbf{0}_{d}-\mathbf{w}_{\star}=-\mathbf{w}_{\star}$.
@@
    This map plays a central role in our analysis and has been studied in similar forms in prior work [23]. Note that ${\mathbf{P}}_{m}$ is an orthogonal projection, which implies that it is symmetric and idempotent. Thus,
@@
    It follows that
    $\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]=Q^{t}\left[{\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]}\right]=Q^{t}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]=Q^{t}\left[{\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)^{\top}}\right]=Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]\,,$
    where $Q^{t}$ denotes $t$ applications of $Q$. The map $Q$ captures the evolution of the error’s second-moment under Kaczmarz updates, offering a tractable approach to analyzing the algorithm’s convergence. The expected loss at step $t$ is given by
    $\mathbb{E}_{\tau}\left[{\mathcal{L}\left({\mathbf{w}_{t}}\right)}\right]$ $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{w}_{t}-\mathbf{y}_{i}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\left({\mathbf{w}_{t}-\mathbf{w}_{\star}}\right)}\right\|^{2}}\right]$
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\left\|{\mathbf{X}\mathbf{z}_{t}}\right\|^{2}}\right]$
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\mathbf{z}_{t}^{\top}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right)}\right]$
@@
    We are now ready to derive the final bound. From Lemma C.7, we have
@@
    Additionally, by Corollary C.5, $Q^{k}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]$ is symmetric and positive semidefinite (PSD). We also note that $\frac{1}{T}\mathbf{X}^{\top}\mathbf{X}$ is symmetric PSD. The key insight from Lemma C.7, combined with the trace product inequality (Lemma C.6), is that it allows the expected loss to be expressed using a polynomial in $Q$. This reformulation simplifies the convergence analysis by reducing it to examining the spectral properties of $Q$. Invoking the trace product inequality, we obtain:
@@
    To clarify, the operator norm of a linear map $H$ is defined as $\left\|{H}\right\|=\sup_{\mathbf{A}\in\mathbb{R}^{d\times d},\left\|{\mathbf{A}}\right\|_{F}=1}\left\|{H\left[{\mathbf{A}}\right]}\right\|_{F}$. The reason for switching from the spectral norm to the Frobenius norm is to enable the use of the spectral mapping theorem to bound the operator norm of $Q^{k-1}\left({I-Q}\right)$, applicable only for inner-product-based norms. We complete the proof by bounding the forgetting using the training loss (Lemma B.1). That is,
@@ C.1 Key Properties and Auxiliary Lemmas
    Let $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, for all $i\in\left[{T}\right]$, $\mathbf{0}\preccurlyeq{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}$. Meaning $Q\left[{\mathbf{A}}\right]$ is PSD as a convex combination of PSD matrices. $\blacksquare$
    Corollary C.4.
@@
    For $n=1$, given by Corollaries C.3 and C.4. For $n>1$, this follows trivially by induction. $\blacksquare$
    Lemma C.6 (Trace Product Inequality).
@@
    Since $\mathbf{0}\preccurlyeq\mathbf{C}=\mathbf{C}^{\top}$, it has a square symmetric PSD root $\mathbf{C}^{1/2}$. Given that $\mathbf{A},\mathbf{B}$ are symmetric and $\mathbf{A}\preccurlyeq\mathbf{B}$, it follows that $\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}\preccurlyeq\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}$ [from 28, Theorem 7.7.2.a]. Applying the cyclic property of the trace and using the fact that for symmetric matrices ordered in the Löwner sense, their traces are also ordered [28, Corollary 7.7.4.d], we obtain
@@
    $\left[\begin{subarray}{c}\text{Im}\left({\mathbf{X}_{i}^{+}\mathbf{X}_{i}}\right)\\ \subseteq\text{Im}\left({\mathbf{X}^{+}\mathbf{X}}\right)\end{subarray}\right]$ $=-\frac{1}{T}\sum_{i=1}^{T}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}+\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{V}_{i}{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}\mathbf{V}_{i}^{\top}.$
@@
    [Corollary C.4]delimited-[]Corollary $\end{subarray}\right]$ $=\operatorname{tr}\left({Q\left[{\mathbf{B}}\right]^{\top}\mathbf{A}}\right)=\operatorname{tr}\left({\mathbf{A}^{\top}Q\left[{\mathbf{B}}\right]}\right)=\langle\mathbf{A},Q\left[{\mathbf{B}}\right]\rangle\,.$
    $\blacksquare$
Proposition C.10.
@@
    From the Rayleigh quotient characterization of eigenvalues, this implies that every eigenvalue $\lambda$ of $Q$ satisfies $0\leq\lambda\leq 1$, i.e., $\sigma(Q)\subset[0,1]\,$. $\blacksquare$
Lemma C.11.
@@
    By Proposition C.9, $Q$ is self adjoint. Thus, we can apply the spectral mapping theorem to the polynomial $x\mapsto x^{n}\left({1-x}\right)$. The eigenvalues of $Q^{n}\left({I-Q}\right)$ are of the form $\lambda^{n}\left({1-\lambda}\right)$, where $\lambda$ is an eigenvalue of $Q$. From Proposition C.10, we know that $\lambda\in\left[0,1\right]$. Using an algebraic property of $\lambda^{n}\left({1-\lambda}\right)$ for $\lambda\in\left[{0,1}\right]$, we conclude that $\lambda^{n}\left({1-\lambda}\right)\in\left[0,\frac{1}{en}\right]$.
    Therefore, $\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en}$. $\blacksquare$
Lemma C.12.
@@
    We first bound $\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$ using the operator norm bound on $Q$ (Proposition C.10):
@@
The proofs in this appendix focus on the properties of forgetting and loss, “translating” them into the language of last-iterate SGD. We then apply our last-iterate results, proved in App. E.
D.1 Proof of Theorem 4.2: A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
  • Recall Theorem 4.2.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 2$ iterations are bounded as,
@@
    Furthermore, Lemma 3.1 states that for all $\mathbf{w}\in\mathbb{R}^{d}$,
@@
    will imply the desired result. Indeed, again by Lemma 3.1, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$. Hence, plugging in $\mathbf{A}=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\Rightarrow\left\|{\mathbf{A}}\right\|=1=\beta$ into Theorem 5.1, SGD with $\eta=1$ guarantees that after $k\geq 1$ gradient steps:
@@
    and therefore $\mathbb{E}\mathcal{L}(\mathbf{w}_{k})\leq\frac{2R^{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}$, which proves the first claim. The second claim follows immediately from Lemma B.1, and we are done. $\blacksquare$
D.2 Proving Theorem 4.3: Main Result for Without Replacement Orderings
  • Recall Theorem 4.3.Under a random ordering without replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\in\{2,\dots,T\}$ iterations are both bounded as,
@@
    From Lemmas B.1 and 3.1, we have
@@
    Combining with Proposition B.2, we get,
@@
Recall that from Eq. 4 in the proof of Lemma B.1, we have
@@
$\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({t}\right)}}\right){\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right)$
$=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)\,.$